# Forecasting the future development in quality and value of professional football players for applications in team management

### v1

@@ Abstract
Transfers in professional football (soccer) are risky investments because of the large transfer fees and high risks involved. Although data-driven models can be used to improve transfer decisions, existing models focus on describing players’ historical progress, leaving their future performance unknown. Moreover, recent developments have called for the use of explainable models combined with uncertainty quantification of predictions. This paper assesses explainable machine learning models based on predictive accuracy and uncertainty quantification methods for the prediction of the future development in quality and transfer value of professional football players. Using a historical data set of data-driven indicators describing player quality and the transfer value of a football player, the models are trained to forecast player quality and player value one year ahead. These two prediction problems demonstrate the efficacy of tree-based models, particularly random forest and XGBoost, in making accurate predictions. In general, the random forest model is found to be the most suitable model because it provides accurate predictions as well as an uncertainty quantification method that naturally arises from the bagging procedure of the random forest model. Additionally, our research shows that the development of player performance contains nonlinear patterns and interactions between variables, and that time series information can provide useful information for the modeling of player performance metrics. Our research provides models to help football clubs make more informed, data-driven transfer decisions by forecasting player quality and transfer value.
@@ 1 Introduction
Transfers in professional football (soccer) are a risky business because the average transfer fee has increased in recent years [1] and because these fees can be characterized as investments with high risks where large fees are involved [2]. Extensive knowledge about players is beneficial to make well-informed decisions about these complex transfer investments in football. This paper shows how football players’ development in quality and monetary value can be predicted to enhance decision making.
Models providing information about player quality and value recently emerged with the evolution of data-driven player performance indicators. Improvements in data-capturing technologies resulted in large data sets containing in-game data about football players, which provide the opportunity to obtain more complex variables on player performance [3, 4]. Numerous player performance indicators have been introduced since then. An example is the expected goals (xG) indicator which values shot chances and shooting ability [5, 6, 7, 8]. Next to such action-specific models, assessment methods exist for general player performance, which can be divided into bottom-up and top-down ratings [9]. Expected threat (xThreat) [10, 11, 12] and VAEP [13, 14, 15, 16] are examples of bottom-up ratings that quantify action quality and use the quality of the actions to create general ratings. Top-down ratings such as plus-minus ratings [17, 18, 19, 20, 21], Elo ratings adjusted for team sports [22], and SciSkill algorithm [23] distribute credit of player performance based on the result of a team as a whole. For the monetary value of players, many models about the estimation of transfer fees and market values have been introduced and provide indicators of the current value of football players [24]. These performance and financial models describe the quality and monetary value of a football player and these can complement traditional scouting reports. This allows managers and technical directors of football clubs to make better-informed transfer decisions.
@@
To this end, two prediction problems are studied: forecasting the quality of a football player and forecasting their transfer value one year ahead. In the first prediction problem, models are trained to predict the development of a top-down quality indicator, the SciSkill [23]. The second prediction problem concerns the prediction of the development of the player value, described by the Estimated Transfer Value (ETV) [25]. The resulting models of these prediction problems offer insight into the question of whether a player will be worth the money in the future. These models thus provide critical insights for the managerial staff of a professional football club.
@@
To further improve the usability of the research results for staff at football organizations, the findings of models should be presented such that they can be utilized by football clubs in practice as stressed by Herold et al. [4]. This means that models should not only be assessed on predictive accuracy but also on explainability and methods for uncertainty quantification [26]. To this end, only explainable supervised learning models are used in this research and the models are assessed on their methods for uncertainty quantification. To focus even more on the applicability for practitioners, the accuracy of the models is partly determined by estimating the loss values of different groups of football players, because certain types of football players are more important from a practitioner’s perspective.
@@
Thus, the aim of this research is to find the most suitable explainable machine learning model to forecast player performance with respect to predictive accuracy and methods for uncertainty quantification. This is done by training supervised models to predict the development of the player quality (SciSkill) and the player value (ETV) one year ahead. To bridge the gap between research and practitioners, only explainable machine learning models are considered and they are assessed on both the predictive quality and the methods for uncertainty quantification. In this way, new models are constructed that can assist in making transfer decisions in football.
@@
This paper is organized as follows: The scientific background on existing literature is given in section 2. The long-term forecasting of player development is then studied for the two prediction problems, of which the methods are described in section 3. The results of the models in the prediction problem of the player quality are presented in section 4 and the results of the player value in section 5. The conclusions are summarized in section 6 followed by a discussion of the research in section 7.
@@ 2.1 Existing indicators
In recent years, the increasing amount of available data in football has driven the introduction of methods to rate individual football players [27]. The performance of a professional football player has traditionally been determined via expert judgment based on video data and statistics describing the frequency of in-game events. Data-driven models have offered the possibility to reduce the bias in assessments and improve the consistency of the judgment of both player quality and player value. In this way, these models have provide new and consistent insights into the quality and monetary value of football players, which has aided in transfer decisions.
@@ 2.1.1 Player performance
Some models define the players’ quality by their actions. Although action-specific models exist [5, 6, 28, 7, 8], methods to assess the quality of all types of actions have been introduced that generalize the action specific models. These general models often define a ‘good’ action as one that increases the probability of scoring and decreases the probability of conceding a goal. The VAEP model by Decroos et al. [13] calculates the probability of scoring given the last three actions including in-game context. Because the involved machine learning techniques are considered a black-box model by practitioners, the authors in [14, 15] introduced methods to make the VAEP model more accessible for practitioners. The research by Decroos and Davis [14] shows that VAEP models with GAMs provide competitive predictive performance while being more interpretable for practitioners. Van Haaren [15] adjusted the VAEP model to use fuzzy assignments to pitch zones instead of a location on the field to relate the model to concepts that are known to practitioners. Further research on the VAEP models has been carried out by Mendes-Neves et al. [16], who have introduced the Intend VAEP (I-VAEP) and Outcome VAEP (O-VAEP) models, which describe the VAEP value of the intent of players and the outcome of their actions, respectively. The authors have studied these VAEP values over time and found that the quality of a football player can be considered as a continuous process over a player’s career. A comparable framework is the xThreat model by Rudd [10], which only considers the current situation, defined by the location of the ball-possessing player, to estimate the probabilities of the ball transitioning to somewhere else on the pitch. This is modeled by a Markov chain, which can be used to describe the probabilities of scoring before and after each action to find the quality of an action. Van Roy et al. [11] have compared the xThreat and VAEP models and have found that, although the xThreat model is more interpretable to practitioners, it can only take into account the position of an action and excludes contextual information such as the position of defenders from its model. This means that there is a trade-off between explainability and the inclusion of in-game context when choosing either VAEP or xThreat models. Van Arem and Bruinsma [12] have extended the xThreat model by including variables describing the defensive situation and height of the ball. This Extended xThreat model can take into account in-game context like the VAEP model, while maintaining explainability as an xThreat model.
@@
The bottom-up ratings like xThreat and VAEP describe the quality of a professional football player by assessing in-game on-the-ball actions. These on-the-ball actions are often offensive actions and they are better at capturing the quality of attackers and attacking midfielders, which indicates that a bias occurs when these models are studied. This bias makes it harder to assess the quality of defensive players and it can be reduced by using top-down models that describe player quality using the lineups and outcomes. Plus-minus ratings are such ratings that were first used in ice hockey and basketball [18], and were later applied to football by Sæbø and Hvattum [17]. For plus-minus ratings, the game is partitioned into game segments that contain the same lineups, which correspond to the data points in the data set. In this data set, the result of the game segment, the goal difference, for example, is the dependent variable. Indicators describing whether a player was active in the segment are the independent variables. Linear regression is then applied to estimate the influence of players on the results. The coefficients of the regression describe the average impact of a player on the game result and give an indicator for player performance over the period of time covered by the data. Because substitutions are infrequent in football, a game does not have many game segments with different lineups. Moreover, football is a low-scoring sport. This creates a situation where a low number of segments with limited distinction in outcomes must be used to infer player quality. To deal with this, other quantities have been used as the dependent variable to describe the result of a segment like the expected number of goals (xG), the expected number of points (xP), and the created VAEP values by Kharrat et al. [18] and Hvattum and Gelade [9]. Pantuso and Hvattum [19] additionally showed the potential of taking age, cards, and home advantage into account, and Hvattum [20] illustrated how separate defensive and offensive ratings can be obtained. Because putting a player in the lineup is an action that a coach performs, De Bacco et al. [21] adapted the plus-minus ratings using a causal model to better describe the influence of the selection of a player. These studies show how plus-minus ratings have been adapted to the application of rating players in football.
@@
Whereas the plus-minus ratings describe the quality of a player using multiple historical games, there also exist models that determine the quality of a player after each game using the lineups and final score. Elo ratings are such ratings and were originally developed to evaluate performance in one-on-one sports. The concept was subsequently adapted to the game of football by Wolf et al. [22]. This adapted algorithm provides ratings for each individual football player and calculates the team rating via the average of players in a game weighted by the number of minutes played. The ratings are then used to predict the match outcome using a fixed logistic function, and after each game, the individual ratings are adjusted. If the outcome is better than predicted, the player rating is increased and if the outcome is worse than expected, it is decreased. The authors also introduced an indicator for player impact to deal with the fact that this rating undervalues good players at below-average teams. The SciSkill [23] is a model that generalizes the Elo rating in several aspects. Instead of only considering one player quality, it describes football players using a defensive rating and an offensive rating. These offensive and defensive scores are then combined to obtain one value, the SciSkill. For each game, the outcome of the game is predicted using a model via an expectation-maximization algorithm. After the prediction, the SciSkill is updated by adjusting the SciSkill values based on the difference with the actual game result. Compared to the Elo algorithm, the SciSkill model extracts more detailed information from the matches and describes the player quality more elaborately.
@@ 2.1.2 Player value
In contrast with player quality, there does exist a ground truth for the concept of player values. The value of a transfer fee is based on the value of the player for each of the involved clubs and historical transfers of similar players [29]. Thus, the historical values of these fees can be used to predict the transfer value of a player albeit at the cost of some selection bias.
@@
A well-known medium that describes the value of a player is Transfermarkt. This company uses crowd estimation to assign values to players [30]. These market values describe the general value of a player and do not take into account the temporary situation of a football player like the current club and contract length. This means that they describe a different quantity than the expected value of a transfer fee. Nonetheless, these values are strongly correlated with the real transfer fees as shown by Herm et al. [30]. Therefore, both market values by Transfermarkt and transfer values are often used interchangeably when describing the monetary value of a football player.
@@
Many studies have been performed to describe the monetary value of a football player based on data. Most of these studies used linear regression as found by Franceschi et al. [24]. These linear models were then used to find what variables have a significant linear dependence on the value of a football player. The study by Franceschi et al. [24] considered 111 trained models that were used in the scientific literature to investigate the transfer value of a football player. The vast majority (85%) of the models are based on ordinary least squares. The authors also have shown the importance of different variables in the considered models. For instance, they have found that age, the square of the age, and the number of matches played by a player are frequently studied variables. These variables are also most often found to be significant. Players frequently increase in value as they get better with experience that is gained over the years, but they also decrease in value as a player loses the potential to improve when getting older. Consequently, the influence of age is often measured with a quadratic term. Similarly, the number of games played can be expected to be an important variable because players gain experience by playing games, which makes them more valuable. Moreover, players who play a lot of games are often the better players on a team. This explains why these variables are important as found by Franceschi et al. [24]. Their study also shows that almost no variables describing defensive behavior are considered when other researchers trained models to describe transfer values. As a consequence, the resulting models can be expected to describe the value of offensive players better than that of defensive players. In addition, the linear models in these studies are mostly trained to determine the influence of variables and the transfer value of a football player. These models are generally not trained and tested for out-of-sample prediction, which limits the application of predicting based on new unseen data.
@@
In contrast, some machine learning studies have been performed to give out-of-sample predictions. This makes it possible to predict the transfer fees for unseen situations or new players. Al-Asadi and Tasdemır [31] have trained multiple models to predict market values of football players with features from the video game FIFA. They have found that a random forest model gives improved prediction over linear methods. Steve Arrul et al. [32] have studied the application of artificial neural networks for the same problem, also considering features from the video game FIFA. They obtain similar loss values as the random forest model of Al-Asadi and Tasdemır [31]. The research by Behravan and Razavi [33] introduces a methodology to train a support vector regression model via particle swarm optimization for the prediction of market values. Although the data set is somewhat similar to the studies above, this model attains worse loss values. In the study of Yang et al. [1], random forests, GAMs, and QAMs are applied to predict the transfer fees of players based on variables describing the player. The authors have inferred from their random forest model that the expenditure of the buying club and the income of the selling club are important features in predicting the transfer fee, as well as the age and the remaining contract duration. This research additionally illustrates how GAMs and QAMs can be used to investigate the dependency between the player transfer value and the given features. The QAM models show that this relation varies for different quantiles, indicating the need to study the influence of models on different groups of players. On the other hand, the GAMs show that the relationship between the transfer values and the features is often nonlinear. A study with extensive types of player performance metrics as features has been performed by McHale and Holmes [2], in which linear regression, linear mixed effects, and XGBoost models have been trained. These models not only include statistics such as the number of minutes played, height, and position, but also plus-minus ratings based on xG, expert ratings from the video game FIFA, and GIM ratings, which are similar to VAEP ratings. The results show that the best predictive performance is attained by the XGBoost model, although the linear mixed effects with the buying and selling clubs as random effects also provide good results. Their results indicate that their model outperforms Transfermarkt market values when predicting the transfer fees on average, although the market values are a better predictor for transfers of more than €20 million. These studies show how supervised learning can be used to obtain models that predict the value of a football player. They show that nonlinear methods generally predict more accurately and that the patterns can differ for different groups of players.

### v2

@@ Abstract
Transfers in professional football (soccer) are risky investments because of the large transfer fees and high risks involved. Although data-driven models can be used to improve transfer decisions, existing models focus on describing players’ historical progress, leaving their future performance unknown. Moreover, recent developments have called for the use of explainable models combined with uncertainty quantification of predictions. This paper assesses explainable machine learning models based on predictive accuracy and uncertainty quantification methods for the prediction of the future development in quality and transfer value of professional football players. The predictive accuracy is studied by training the models to predict the quality and value of players one year ahead. This is carried out by training them on two data sets containing data-driven indicators describing the player quality and player value in historical settings. In general, the random forest model is found to be the most suitable model because it provides accurate predictions as well as an uncertainty quantification method that naturally arises from the bagging procedure of the random forest model. Additionally, this research shows that the development of player performance contains nonlinear patterns and interactions between variables, and that time series information can provide useful information for the modeling of player performance metrics. The resulting models can help football clubs make more informed, data-driven transfer decisions by forecasting player quality and transfer value.
@@ 1 Introduction
Transfers in professional football (soccer) are a risky business because the average transfer fee has increased in recent years [1] and because these fees can be characterized as investments with high risks where large fees are involved [2]. Extensive knowledge about players is beneficial in making well-informed decisions about these complex transfer investments in football. This paper shows how football players’ development in quality and monetary value can be predicted to enhance decision-making.
Models providing information about player quality and value have recently emerged with the evolution of data-driven player performance indicators. Improvements in data-capturing technologies resulted in large data sets containing in-game data about football players, which provide the opportunity to obtain more complex variables on player performance [3, 4]. Numerous player performance indicators have been introduced since then. An example is the expected goals (xG) indicator, which values shot chances and shooting ability [5, 6, 7, 8]. Next to such action-specific models, assessment methods exist for general player performance, which can be divided into bottom-up and top-down ratings [9]. Expected threat (xThreat) [10, 11, 12] and VAEP [13, 14, 15, 16] are examples of bottom-up ratings that quantify action quality and use the quality of the actions to create general ratings. Top-down ratings such as plus-minus ratings [17, 18, 19, 20, 21], Elo ratings adjusted for team sports [22], and SciSkill algorithm [23] distribute credit of player performance based on the result of a team as a whole. For the monetary value of players, many models about the estimation of transfer fees and market values have been introduced and provide indicators of the current value of football players [24]. These performance and financial models describe the quality and monetary value of a football player, and these can complement traditional scouting reports. This allows managers and technical directors of football clubs to make better-informed transfer decisions.
@@
To this end, two prediction problems are studied: forecasting the quality of a football player and forecasting their transfer value one year ahead. In the first prediction problem, models are trained to predict the development of a top-down quality indicator, the SciSkill [23]. The second prediction problem concerns the prediction of the development of the player value, described by the Estimated Transfer Value (ETV) [25]. A forecasting horizon of one year is selected in this study as it aligns with the length of one full season in football. The resulting models of these prediction problems offer insight into the question of whether a player will be worth the money in the future. These models thus provide critical insights for the managerial staff of a professional football club.
@@
To further improve the usability of the research results for staff at football organizations, the findings of models should be presented such that they can be utilized by football clubs in practice, as stressed by Herold et al. [4]. This means that models should not only be assessed on predictive accuracy but also on explainability and methods for uncertainty quantification [26]. To this end, only explainable supervised learning models are used in this research, and the models are assessed on their methods for uncertainty quantification. To focus even more on the applicability for practitioners, the accuracy of the models is partly determined by estimating the loss values of different groups of football players, because certain types of football players are more important from a practitioner’s perspective.
@@
The aim of this research is to find the most suitable explainable machine learning model to forecast player performance with respect to predictive accuracy and methods for uncertainty quantification. To find this, several explainable supervised learning models are studied. Their methods for uncertainty quantification are studied based on the literature. The predictive performance is assessed by implementing them to forecast the player quality (SciSkill) and the player value (Estimated Transfer Value) of football players one year ahead. The predictive quality is studied on both the general population of players and on subgroups of players that are more interesting for practitioners, such as young or high-value players. An overview of this process is visualized in Figure 1. These results will then be combined to find an answer to the research question.
Figure 1: A visualization of the structure of this paper.
The main contribution of this paper is the illustration of how extra value can be added to football player KPIs by forecasting future values. The results of this study provide knowledge on what models are most suitable for application by practitioners, lowering the threshold for real-life implementation. By taking explainability, uncertainty quantification, and performance on important subpopulations of players into account, this research is presented such that the results can be used easily by practitioners. In this way, this paper also contributes to bridging the gap between academic research and practitioners.
@@
This paper is organized as follows: The scientific background on existing literature is given in Section 2. The long-term forecasting of player development is then studied for the two prediction problems, for which the methods are described in Section 3. The results of the models in the prediction problems of the player quality and player value are presented in Section 4. The conclusions of this research are then summarized in Section 5, followed by a discussion of the implications and future directions in Section 6.
@@ 2 Background
First, the considered models are discussed along with an assessment of their methods for uncertainty quantification. Then, more background is given on existing models for football player quality and value.
2.1 Supervised models
Supervised learning models provide a possibility to forecast the player performance indicators in the future. Based on input variables, they predict output variables, and they can be trained based on an existing data set for prediction problems. Such a data set $\mathcal{D_{\text{train}}}=\{(X_{1},y_{1}),\ldots,(X_{n},y_{n})\}$ consists of features $X_{i}$ describing characteristics about a data point and labels $y_{i}$ describing the actual values. In this research, the features $X_{i}$ are values that describe the situation of a player at a specific point in time, and the labels are the player’s performance indicator exactly one year later. By constructing these features and labels for multiple points through time and for different players, a data set can be created to train the models.
2.1.1 Considered models
To align the research with practitioner needs, explainable models are studied in this paper. The considered models can be divided into linear models, tree-based models, and kNN-based models, and they will now be shortly described.
The first linear model considered is ordinary least squares (OLS), sometimes called (multiple) linear regression as described by Hastie et al. [27, Section 3.2]. This model assumes that the true relation is linear $y_{i}=X_{i}\beta+\varepsilon_{i}$, where $\varepsilon_{i}$ is normally distributed noise. OLS minimizes the residual sum of squares
$RSS(\beta)=(y-X\beta)^{T}(y-X\beta),$
and has the lowest mean square error of all unbiased estimators. Lasso regression [28] is another linear model that introduces a bias in the model by minimizing the penalized residual sum of squares $RRS(\beta)+\lambda\sum_{i=1}^{p}|\beta_{i}|$. By introducing this bias, it is possible to provide more accurate predictions. The third linear model is the linear mixed effects (LME) model, which assumes that
$y_{i}|b_{i}=X_{i}\beta+Z_{i}b_{i}+\varepsilon_{i},$
where $\varepsilon_{i}\sim N(0,\sigma^{2}\Lambda_{i})$[29]. The random variable $Z_{i}$ is called the random effect of group $i$ and can be used to model the influence of a specific random attribute that should not influence the predictions, e.g., nationality. Because of their linearity, these models are explainable and suitable for applications.
The CART decision tree [27, Section 9.2] is a model that recursively splits the feature space by taking splits of the form $\{X|X_{j}\leq s\},\{X|X_{j}>s\}$ such that it minimizes the sum of squares within each split. As described by Hastie et al. [27, p.312], decision trees generally suffer from a high variance. A random forest model solves this by repeatedly training decision trees on resampled data [27, Chapter 15]. The random forest model then predicts by taking the average prediction of these decision trees. Boosting is another way of improving decision trees. A boosting algorithm fits decision trees sequentially and reweighs the data points for which the model has a bad performance [27, Chapter 10]. The XGBoost [30] combines several techniques, such as regularization and feature quantile estimates, to provide such a boosting algorithm. Generally, the splitting behavior in tree-based algorithms mimics the if-then reasoning of humans. Additionally, they can provide feature importances by considering how much a specific feature improves the predictions on the training set. This makes the tree-based models explainable and thus suitable models.
The last type of model considered in this paper is the k-nearest neighbor (kNN) model. Given the features, these models search for the most similar data points in the training data set and predict the associated $y_{i}$-value by taking the average of the $k$-nearest neighbors [27, Subsection 2.3.2]. The closest neighbors can be weighted more heavily as illustrated by Dudani [31]. The ReliefF algorithm by [32] calculates feature importances by looking at the probability of having different values considering the neighbors of a data point. A downside of the kNN models is that they suffer from the curse of dimensionality as illustrated in [27, Section 2.5]. Nonetheless, the kNN models can provide examples of training data points to explain predictions. In the context of the current study, this would be players who have been in similar situations in history, which gives a very intuitive interpretation to practitioners.
2.1.2 Uncertainty quantification
To better align research with application in practice, the models in this study are also assessed on their methods for uncertainty quantification. Although quantile regression could be applied to obtain a form of uncertainty quantification, this requires an extra model. This makes quantile regression less useful for applications in practice. Therefore, models with uncertainty quantification methods that do not need to alter the model or results are favorable.
Linear regression has an underlying theory that gives prediction intervals as described by Neter et al. [33]. These prediction intervals are based on assumptions that are unlikely to hold for the prediction problems in this paper, such as the normality of errors. Nonetheless, they do give an indication of the uncertainty of the prediction. Similarly, the kNN models provide the neighbors, which are a group of similar data points. Uncertainty quantification can be obtained by taking the minimal and maximal values of the dependent values within these neighbors if the number of neighbors $k$ is of significant size. Lastly, the bagging procedure of the random forest model can be utilized to obtain uncertainty quantification for the predictions, as described by Wager et al. [34]. This method utilizes the different decision trees in the random forest to quantify the uncertainty in the prediction. From these properties, it is concluded that the linear regression, random forest, and kNN-based models are favorable with respect to uncertainty quantification.
@@ 2.2 Existing indicators
In recent years, the increasing amount of available data in football has driven the introduction of methods to rate individual football players [35]. The performance of a professional football player has traditionally been determined via expert judgment based on video data and statistics describing the frequency of in-game events. Data-driven models have offered the possibility to reduce the bias in assessments and improve the consistency of the judgment of both player quality and player value, as discussed below. In this way, these models have provided new and consistent insights into the quality and monetary value of football players, which has aided in transfer decisions.
@@ 2.2.1 Player performance
Some models define the players’ quality by their actions. Although action-specific models exist [5, 6, 36, 7, 8], methods to assess the quality of all types of actions have been introduced that generalize the action-specific models. These general models often define a ‘good’ action as one that increases the probability of scoring and decreases the probability of conceding a goal. The VAEP model by Decroos et al. [13] calculates the probability of scoring given the last three actions, including in-game context. Because the involved machine learning techniques are considered a black-box model by practitioners, the authors in [14, 15] introduced methods to make the VAEP model more accessible for practitioners. A comparable, more interpretable framework is the xThreat model by Rudd [10], which only considers the current situation, defined by the location of the ball-possessing player, to estimate the probabilities of the ball transitioning to somewhere else on the pitch. This is modeled by a Markov chain, which can be used to describe the probabilities of scoring before and after each action to find the quality of an action. Van Roy et al. [11] have compared the xThreat and VAEP models and have found that, although the xThreat model is more interpretable to practitioners, it can only take into account the position of an action and excludes contextual information such as the position of defenders from its model. This means that there is a trade-off between explainability and the inclusion of in-game context when choosing either VAEP or xThreat models. Van Arem and Bruinsma [12] have extended the xThreat model by including variables describing the defensive situation and height of the ball. This Extended xThreat model can take into account in-game context like the VAEP model, while maintaining explainability as an xThreat model.
@@
The bottom-up ratings like xThreat and VAEP describe the quality of a professional football player by assessing in-game on-the-ball actions. These on-the-ball actions are often offensive actions, and they are better at capturing the quality of attackers and attacking midfielders, which indicates that a bias occurs when these models are studied. This bias makes it harder to assess the quality of defensive players, and it can be reduced by using top-down models that describe player quality using the lineups and outcomes. Plus-minus ratings are such ratings that were first used in ice hockey and basketball [18], and were later applied to football by Sæbø and Hvattum [17]. For plus-minus ratings, the game is partitioned into game segments that contain the same lineups, which correspond to the data points in the data set. In this data set, the result of the game segment, the goal difference, for example, is the dependent variable. Indicators describing whether a player was active in the segment are the independent variables. Linear regression is then applied to estimate the influence of players on the results. The coefficients of the regression describe the average impact of a player on the game result and give an indicator for player performance over the period of time covered by the data. Because substitutions are infrequent in football, a game does not have many game segments with different lineups. Moreover, football is a low-scoring sport. This creates a situation where a low number of segments with limited distinction in outcomes must be used to infer player quality. To deal with this, other quantities have been used as the dependent variable to describe the result of a segment, like the expected number of goals (xG), the expected number of points (xP), and the created VAEP values by Kharrat et al. [18] and Hvattum and Gelade [9]. Pantuso and Hvattum [19] additionally showed the potential of taking age, cards, and home advantage into account, and Hvattum [20] illustrated how separate defensive and offensive ratings can be obtained. Because putting a player in the lineup is an action that a coach performs, De Bacco et al. [21] adapted the plus-minus ratings using a causal model to better describe the influence of the selection of a player. These studies show how plus-minus ratings have been adapted to the application of rating players in football.
@@
Whereas the plus-minus ratings describe the quality of a player using multiple historical games, there also exist models that determine the quality of a player after each game using the lineups and final score. Elo ratings are such ratings and were originally developed to evaluate performance in one-on-one sports. The concept was subsequently adapted to the game of football by Wolf et al. [22]. This adapted algorithm provides ratings for each individual football player and calculates the team rating via the average of players in a game weighted by the number of minutes played. The ratings are then used to predict the match outcome using a fixed logistic function, and after each game, the individual ratings are adjusted. If the outcome is better than predicted, the player rating is increased, and if the outcome is worse than expected, it is decreased. The authors also introduced an indicator for player impact to deal with the fact that this rating undervalues good players at below-average teams. The SciSkill [23] is a model that generalizes the Elo rating in several aspects. Instead of only considering one player’s quality, it describes football players using a defensive rating and an offensive rating. These offensive and defensive scores are then combined to obtain one value, the SciSkill. For each game, the outcome of the game is predicted using a model via an expectation-maximization algorithm. After the prediction, the SciSkill is updated by adjusting the SciSkill values based on the difference with the actual game result. Compared to the Elo algorithm, the SciSkill model extracts more detailed information from the matches and describes the player quality more elaborately.
@@ 2.2.2 Player value
In contrast with player quality, there does exist a ground truth for the concept of player values. The value of a transfer fee is based on the value of the player for each of the involved clubs and historical transfers of similar players [37]. Thus, the historical values of these fees can be used to predict the transfer value of a player, albeit at the cost of some selection bias.
@@
A well-known medium that describes the value of a player is Transfermarkt. This company uses crowd estimation to assign values to players [38]. These market values describe the general value of a player and do not take into account the temporary situation of a football player, like the current club and contract length. This means that they describe a different quantity than the expected value of a transfer fee. Nonetheless, these values are strongly correlated with the real transfer fees as shown by Herm et al. [38]. Therefore, both market values by Transfermarkt and transfer values are often used interchangeably when describing the monetary value of a football player.
@@
The literature review by Franceschi et al. [24] showed that many studies have been performed to describe the monetary value of a football player based on data. The authors found that most of these studies used linear regression to find which variables have a significant linear dependence on the value of a football player. The review by Franceschi et al. [24] considered 111 trained models that were used in the scientific literature to investigate the transfer value of a football player. The vast majority (85%) of the models are based on ordinary least squares. The authors have also shown the importance of different variables in the considered models. For instance, they have found that age, the square of the age, and the number of matches played by a player are frequently studied variables. These variables are also most often found to be significant. Players frequently increase in value as they get better with experience that is gained over the years, but they also decrease in value as a player loses the potential to improve when getting older. Consequently, the influence of age is often measured with a quadratic term. Similarly, the number of games played can be expected to be an important variable because players gain experience by playing games, which makes them more valuable. Moreover, players who play a lot of games are often the better players on a team. This explains why these variables are important, as found by Franceschi et al. [24]. Their study also shows that almost no variables describing defensive behavior are considered when other researchers trained models to describe transfer values. As a consequence, the resulting models can be expected to describe the value of offensive players better than that of defensive players. In addition, the linear models in these studies are mostly trained to determine the influence of variables on the transfer value of a football player, and they are generally not trained and tested for out-of-sample prediction, which limits the application of predicting based on new, unseen data.
@@
In contrast, Al-Asadi and Tasdemır [39] have trained multiple models to predict market values of football players with features from the video game FIFA. They have found that a random forest model gives improved predictions over linear methods. Steve Arrul et al. [40] have studied the application of artificial neural networks for the same problem, also considering features from the video game FIFA. They obtain similar loss values as the random forest model of Al-Asadi and Tasdemır [39]. The research by Behravan and Razavi [41] introduces a methodology to train a support vector regression model via particle swarm optimization for the prediction of market values. Although the data set is somewhat similar to the studies above, this model attains worse loss values. In the study of Yang et al. [1], random forests, GAMs, and QAMs are applied to predict the transfer fees of players based on variables describing the player. The authors have inferred from their random forest model that the expenditure of the buying club and the income of the selling club are important features in predicting the transfer fee, as well as the age and the remaining contract duration. This research additionally illustrates how GAMs and QAMs can be used to investigate the dependency between the player transfer value and the given features. The QAM models show that this relation varies for different quantiles, indicating the need to study the influence of models on different groups of players. On the other hand, the GAMs show that the relationship between the transfer values and the features is often nonlinear. A study with extensive types of player performance metrics as features has been performed by McHale and Holmes [2], in which linear regression, linear mixed effects, and XGBoost models have been trained. These models not only include statistics such as the number of minutes played, height, and position, but also plus-minus ratings based on xG, expert ratings from the video game FIFA, and GIM ratings, which are similar to VAEP ratings. The results show that the best predictive performance is attained by the XGBoost model, although the linear mixed effects with the buying and selling clubs as random effects also provide good results. Their results indicate that their model outperforms Transfermarkt market values when predicting the transfer fees on average, although the market values are a better predictor for transfers of more than €20 million. These studies show how supervised learning can be used to obtain models that predict the value of a football player. They show that nonlinear methods generally predict more accurately and that the patterns can differ for different groups of players.